If the amplitude of z–2–3i is pi/4 then the locus of z=x+iyis

Grade:12

1 Answers

Ritesh Khatri
6 years ago
let us consider points z1 = 2+3i and z = x+iy
now z – zrepresents vector joining z and z1 whose slope is tan(pi/4)
since  slope is constant z can only move along straight line hence locus is a straight line